The Star One Bus service is considering raising price of its tickets from current fare of

Rupees 300 per seat.

 

The Demand schedule for it’s regular route is:

Price (Rupees per ticket)

Quantity Demanded (tickets per month)

200

900

300

750

400

600

500

450

600

300

 

3a. Calculate the price elasticity of demand if Star One raises the fare from Rs.300 to Rs.400. Is the demand price elastic or inelastic for this fare rise? 


1
Expert's answer
2021-04-07T07:25:24-0400

Solution:

3a.). The price elasticity of demand ="\frac{percentage \;change\; in \;quantity\; demanded}{percentage\; change\; in \;price}"


% change in quantity = "\frac{Q_{2} - Q_{1} }{(Q_{2} + Q_{1}) \div2}"

Q1 = 750

Q2 = 600

 

= "\frac{600 - 750 }{(600 + 750) \div2} \times 100" = "\frac{-150}{675} \times 100 = -0.2222 \times 100 = - 22.22\%"


% change in price = "\frac{P_{2} - P_{1} }{(P_{2} + P_{1}) \div2} \times 100"

P1 = 300

P2 = 400


= "\frac{400 - 300 }{(400 + 300) \div2} \times 100 = \frac{100}{350} \times 100 = -0.29 \times 100 = -29\%"


PED ="\frac{percentage \;change\; in \;quantity\; demanded}{percentage\; change\; in \;price}"


 ="\frac{-22.22\%}{29\%} = -0.76"


PED = 0.76

The demand for this fare rise is inelastic since it is less than one.

It means that a change in price causes a smaller percentage change in demand, which suggests that the service has fewer substitutes.